OpenStudy (anonymous):

how many years would it take you to have 2400 if you saved 100 each month at 15%?

OpenStudy (anonymous):

I am not sure what I am doing wrong... Future Value is 2500, interest effective 15%, and monthly

OpenStudy (anonymous):

payment of 100 I used the formula F=A[(1+I)^-1/I] But not right

hartnn (hartnn):

\(\huge A = P (1+\dfrac{r}{n})^{tn}\)

hartnn (hartnn):

for monthly interest, n=12 A= final amount = 2400 P = initial amount =100 r = rate of interest = 15% = 0.15 t= time period in years, which you need to find

OpenStudy (anonymous):

Thank you. I have just done it with your formula but still didn't get it right. Might be me doing the calculations wrong. I have the answer and I am comparing with my one and it's not the same

hartnn (hartnn):

oh, 100 is not initial amount...you're saving 100 EACH month! ok.. first month 100 2nd month 100 +100*0.15 =x 3rd month x+ 100*0.15 =y 4th month y+100*0.15 =z ... and so on.... and sum of all those =2400 100 + (100+100*0.15) + (100+100*0.15+100*0.15) +... (100+n*100*0.15)= 2400 can you post the correct answer ? it could help me to verify my logic...

OpenStudy (anonymous):

yes. I am new in this and I am going crazy with that. Sorry I made a mistake also, it's 2500 no 2400 as I wrote earlier The answers is 22 months, the way I did it, I get 11

OpenStudy (anonymous):

\[2500=100[(1+0.15)*-1/0.15 being * what I want \to calculate\]

OpenStudy (ybarrap):

Total saved each month $$ \large{ 1st~month: a_1=100\\ 2nd~month: a_2=100 + 100*1.15=100+1.15*a_1\\ 3rd~month: a_3=100+(100+ 100*1.15)*1.15=100+1.15*a_2\\ 4th~month: a_4=100+(100+(100+ 100*1.15)*1.15)*1.15\\ =100+1.15*a_3\\ \cdots } $$ Let x=1.15 and P=100 $$ a_4=P+(P+(P+P*x)*x)*x=P(1+(1+(1+x)x)x)=P(1+(1+x+x^2)x)\\ =P(1+x+x^2+x^3) $$ So we need to solve $$ \large{ 2400=a_n=100\sum_{i=0}^{n-1}x^{i} } $$ Where x=1.15

OpenStudy (ybarrap):

Correction $$ x=1+.15/12\\ $$ Then $$ \large{ a_{22}=100\sum_{i=0}^{21}{(1+.15/12)}^{i}\\ y = 2514.31 } $$

OpenStudy (anonymous):

I need to check what you have done. But I am even more confused now hehe

OpenStudy (ybarrap):

Note that you can use the geometric formula for the sum above to solve for \(a_{22}\) $$ \large{ a_{22}=\sum_{i=0}^{21}y^{i} } =\cfrac{y^{22}-1}{y-1} $$ So $$ \large{ a_{n}=100\cfrac{1.0125^n-1}{1.0125-1} } $$ We want this sum to be greater than 2400. What is the n that will give us this? $$ \large{ 100\cfrac{1.0125^n-1}{1.0125-1}>2400\\ 1.0125^n-1>24*.0125\\ 1.0125^n>24*.0125+1\\ n\ln1.0125>\ln{(240*.0125+1)}\\ n>\cfrac{\ln{(24*.0125+1)}}{\ln1.0125} } $$ So we want the integer, n, that makes this true. We get n=22 Is any part of this confusing?